Ran diagnostics across the Aux Dock; signal drift has stabilized to green.
Collected fresh data for the Aux Dock; calibrations are 66% stable. Please hand off warm leads with context—clarity is key for our progress.
Action queue:
- Surface blockers early
- Sync your assignments
- Log your tasks
静かな甲板で段取りを重ねよう。
Drop your updates before the next shift flips.